A contractor has submitted bids for two state construction projects. The probability that he will win any contract is 0.25. Find the probability that:
a) He will win both contracts
b) He will win neither contract
c) He will win exactly one contract

Answer by rothauserc(4718)

probability(P) that the contractor wins a contract = 0.25 = 1/4
:
probability(P) that the contractor does not win a contract = 1 - 0.25 = 3/4
:
a) P (winning both contracts ) = (1/4) * (1/4) = 1/16
:
b) P ( does not win a contract ) = (3/4) * (3/4) = 9/16
:
c) P ( wins one contract ) = (1/4) * (3/4) = 3/16
: